OpenStudy (anonymous):

Medal for best answer 4n=1/64

hartnn (hartnn):

is it \(\Large 4^n = \dfrac{1}{64}\) ?

geerky42 (geerky42):

IS IT?!?!?!

OpenStudy (milkncookies):

I think he meant for you just to solve for N....

geerky42 (geerky42):

Yeah, but we can't tell if it's \(4^n = \dfrac{1}{64}\) or \(4n = \dfrac{1}{64}\)

OpenStudy (milkncookies):

What number times 4 = 1/64?

OpenStudy (milkncookies):

Oh...

OpenStudy (milkncookies):

Pretty sure it's 4n...

geerky42 (geerky42):

yeah, but you know some of users here forgot or don't know "^" and in your case, answer will be a little ugly so that's why we assumed he meant to say 4^n

geerky42 (geerky42):

@noneyabusiness please say something...

OpenStudy (milkncookies):

Maybe he left?

OpenStudy (anonymous):

sorry had to do something...

geerky42 (geerky42):

ok, we just want to make sure, do you mean 4n or 4^n?

OpenStudy (anonymous):

\[4n\]=1/64

OpenStudy (anonymous):

I am sorry for leaving and I am a girl

geerky42 (geerky42):

ok, in case, just divide both side by 4. do you need help with that?

OpenStudy (anonymous):

ok so I divide 1/64 by 4? and get... 1/256

OpenStudy (anonymous):

lol cool names

geerky42 (geerky42):

Yeah, that it.

geerky42 (geerky42):

n = 1/256

OpenStudy (anonymous):

YAY THANK YOU!!!

geerky42 (geerky42):

when you need to find x or anything, all you have to do is to "isolate" it. You get 4 to other side.

geerky42 (geerky42):

you are welcome.

OpenStudy (anonymous):

Okeedokee